OpenStudy (anonymous):

Which is the equation of the line passing through (0, 8) and (4, 0)?

OpenStudy (anonymous):

y=-2*x+8 you can get the slope as \[\frac{ 0-8 }{ 4-0 }\] and the y-cut is 8, since it is the yvalue for x=0

OpenStudy (anonymous):

Given the linear equation x + y = 5, find the slope of its graph. –1 1 5

OpenStudy (anonymous):

can u help with these